# Quantum Polar Metric Learning: Efficient Classically Learned Quantum Embeddings

### 11 Jan 2024 v2

@@ 1. Introduction
We find that QPMeL outperforms the previous QMLeL method both in capability and computational complexity. We achieve a high degree of separation between all 10 classes using only 3 qubits and 4 gates per qubit. This is a 3x reduction compared to the 2-layer QAOA implementation of (Lloyd et al., 2020). We also demonstrate that QPMeL outperforms classical networks with similar structures to our classical head, presenting a promising avenue for future research on fully classical models with quantum loss functions.
@@ 2.1.1. Learned Encoding Schemes
A proposed solution to this problem was in the form of Learned Encoding Schemes for QML tasks. In this paradigm, a classical computer learns a lower dimensional representation of the data via methods such as PCA or Deep Neural Networks(DNNs) and these are then used as the input to a QML circuit. However, as shown by Lloyd et al (Lloyd et al., 2020), the main challenge with this approach is the poor utilization of the Hilbert Space as the classical model creates real-valued nonperiodic bottlenecks that do not translate well to the complex, exponential and periodic natural of Hilbert Space.
@@ 4.1.1. CNN Backbone
The classical head uses convolution blocks consisting of CONV + ReLU + MaxPool layers, a dense block with 3 Dense + GeLU layers with reducing dimensionality. GeLU has better convergence properties in Dense layers and hence is chosen over ReLU. The architecture can be seen in 2.
@@ 4.2.1. Encoding Circuit
The encoding circuit is used to create the state from the classical embeddings. The structure (as shown in Fig2(a)) consists of and gates separated by a layer of cyclic gates for entanglement. Our experiments show that this structure performs similarly to or slightly better than an structure which more naturally shows polar learning. The choice of is motivated by the variable entanglement property as also observed by (Lloyd et al., 2020).
@@ 4.2.2. Embedding State and Learnable Parameters
The final state produced by our Encoding circuit as seen in Fig 2(a) would be:
@@ 4.2.3. Training Circuit
QPMeL uses separate circuits for training and inference, with 2 main differences - (1). The SWAP test extension requires 2 copies of the encoding circuit (2). Residual Corrections that are only used in the training process. We can see the green regions in Fig 2(b) use our encoding circuit but add the QRC parameters (,) to the rotations.
@@ 4.3. Quantum Residual Corrections (QRC)
To overcome this issue we propose a novel new method we name ”QRC”. We add learnable parameters and to the angles of the and gates respectively as seen in Fig.3. Due to their shallowness and input independence, they are affected by smaller gradients faster and act as ”noise barrier” allowing our classical to learn faster. During inference, we add these weights to our classical model as seen in Fig 2
@@ 4.4. Fidelity Triplet Loss
A key difference from the classical counterpart is using rather than . This is a natural result of the difference between Fidelity and MSE distance metrics. The classical formulation tries to minimize AP Distance, as for similar features. However, in the quantum case, we try to maximize the fidelity, as for similar states.
@@ 5. Experimental Set-Up
5.0.1. MinMax Metric:
We propose the ”MinMax” metric to quantify our embedding performance. The metric is feature space agnostic as it directly operates on the pre-computed distances using the native space distance metric (MSE or fidelity). However, as the metric trends are reversed for MSE and Fidelity, we present the results separately and use for Fidelity. The metric computes the distance of the and , which corresponds to the worst case for both and represents it as a % . These are computed on the averages seen in the heatmap to avoid outliers.
@@ 5.0.1. MinMax Metric:
(Decision Boundry at 0.1)
(No Decision Boundry)
(No Decision Boundry)
(Decision Boundry at 1.5)
6. Empirical Results and Observations
6.0.1. Classical Baselines:
6.0.2. Quantum Baselines:
(Perfect Separation)
(Smaller Perfect Separation)
(Multiple Outliers)
(4-6 pair is not separable)
Hilbert Compare.
6.1. 3x better separation using 1/2 the number of gates and circuit depth compared to QMeL
The original QMeL framework (Lloyd et al., 2020) as explained in Sec. 3 used an overlap loss that did not scale well to multi-class tasks. The modified version with our Fidelity Triplet Loss also failed to create a decision boundary as seen in Table.1 where it produces a -ve difference. When we apply a more robust separation method for the pre-quantum step (QMeL+) the difference in Table.1 is +ve implying that a decision boundary can be made but the magnitude is only 24% of implying that the clusters are close together.
In contrast, QPMeL produces a +ve difference that is 91% of , showing 3x improvement over QMeL+. We see this trend reflected in Fig.4(a) and Fig.4(b) with the difference in magnitude between the diagonal and everywhere else. Additionally, looking at the parameters in Table 2 QPMeL utilizing 1/2 (9 vs 21) the number of gates, 1/2 (5 vs 11) the circuit depth and (11k vs 16k)20% fewer classical parameters compared to QMeL/QMeL+
@@ 6.2. Quantum loss functions may improve the learning of classical models
Look at Table 3, both sigmoid models produce -ve differences which indicate the absence of a decision boundary. The QPMeL classical head has an identical structure to the Scaled Sigmoid model with only a difference in the loss function. However, it produces a +ve difference of 88% of , implying that a strong decision boundary exists.
@@ 6.3. QPMeL strongly Couples Euclidean and Hilbert embedding seperation
Despite never being trained on any Euclidean loss function such as MSE-Triplet Loss, the QPMeL classical head produces a strong separation between classes as seen by the 88% of in Table 3. We can also see in Fig 3(a) that this separation is well distributed across all pairs.
@@ 6.4. QRC enables the classical head to learn decision boundaries in Hilbert Space
From the results in Table. 1, we can see that the No Residual model fails to learn a clear decision boundary. However, looking Fig 4(c), while the model fails to create a decision boundary, the issues are localized to specific pairs (ex. (8,2)), while all other regions remain well separated. We believe this is due to smaller gradients which require more training time and data to learn these corner cases. However, PQCs are known to have barren plateaus which can make training unstable.
In contrast, our model can learn a decision boundary for all classes. Fig 4(d), shows that even before applying our corrections the encoder struggles with a single pair. This implies that QRC is helping the classical model learn the corner cases faster. This is further supported by the fact that the QRC is only used during training and not during inference. This hints that the residual framework eases the task on the classical head allowing it to learn faster and more robustly.

### 27 Feb 2024 v3

@@ 1. Introduction
We find that QPMeL outperforms the previous QMeL method both in capability and computational complexity. We achieve a high degree of separation between all 10 classes using only 3 qubits and 4 gates per qubit. This is a 3x reduction compared to the 2-layer QAOA implementation of (Lloyd et al., 2020). We also demonstrate that QPMeL outperforms classical networks with similar structures to our classical head, presenting a promising avenue for future research on fully classical models with quantum loss functions.
@@ 2.1.1. Learned Encoding Schemes
A proposed solution to this problem was in the form of Learned Encoding Schemes for QML tasks. In this paradigm, a classical computer learns a lower dimensional representation of the data via methods such as Principle Component Analysis (PCA) or Deep Neural Networks(DNNs) and these are then used as the input to a QML circuit. However, as shown by Lloyd et al (Lloyd et al., 2020), the main challenge with this approach is the poor utilization of the Hilbert Space as the classical model creates real-valued nonperiodic bottlenecks that do not translate well to the complex, exponential and periodic natural of Hilbert Space.
@@ 4.1.1. CNN Backbone
The classical head uses convolution blocks consisting of CONV + ReLU + MaxPool layers, a dense block with 3 Dense + GeLU layers with reducing dimensionality. GeLU has better convergence properties in Dense layers and hence is chosen over ReLU. The architecture can be seen in Fig. 2.
@@ 4.2.1. Encoding Circuit
The encoding circuit is used to create the state from the classical embeddings. The structure (as shown in Fig. 2(a)) consists of and gates separated by a layer of cyclic gates for entanglement. Our experiments show that this structure performs similarly to or slightly better than an structure which more naturally shows polar learning. The choice of is motivated by the variable entanglement property as also observed by (Lloyd et al., 2020).
@@ 4.2.2. Embedding State and Learnable Parameters
The final state produced by our Encoding circuit as seen in Fig. 2(a) would be:
@@ 4.2.3. Training Circuit
QPMeL uses separate circuits for training and inference, with 2 main differences - (1). The SWAP test extension requires 2 copies of the encoding circuit (2). Residual Corrections that are only used in the training process. We can see the green regions in Fig. 2(b) use our encoding circuit but add the QRC parameters (,) to the rotations.
@@ 4.3. Quantum Residual Corrections (QRC)
To overcome this issue we propose a novel new method we name ”QRC”. We add learnable parameters and to the angles of the and gates respectively as seen in Fig.3. Due to their shallowness and input independence, they are affected by smaller gradients faster and act as ”noise barrier” allowing our classical to learn faster. During inference, we add these weights to our classical model as seen in Fig.2
@@ 4.4. Fidelity Triplet Loss
A key difference from the classical counterpart is using rather than . This is a natural result of the difference between Fidelity and Mean Squared Error(MSE) distance metrics. The classical formulation tries to minimize AP Distance, as for similar features. However, in the quantum case, we try to maximize the fidelity, as for similar states.
5. Analysis
5.1. QPMeL as Dense Angle Encoding
Dense angle encoding is a variation of angle encoding that exploits relative phase as a degree of freedom to encode 2 real features per qubit. Another way for us to frame this would be to consider it a method to encode 2-D data points onto a single qubit as shown by (LaRose and Coyle, 2020). The structure of Dense Angle Encoding is shown below in Eq. 3.
A key point to note is that dense angle encoding essentially encodes data onto to the polar coordinates of a single qubit. This is apparent when comparing the structure of the polar representation of a single qubit state to the dense angle encoding representation. I can trivially prove that QPMeL for single qubits is equivalent to Dense Angle Encoding and hence to Polar Encoding.
Taking our single qubit state as,
Without loss of generality, we can consider the global phase of the system as , which would reduce our initial mapping to :
Eq. 4 is identical to the polar representation of a single qubit state.
Recall that and are scaled by before encoding into the qubit. Therefore, replacing them in Eq. 4 we get:
Where, are the weights from our classical network that yield and respectively. is the sigmoid function and is the image. Taking and , we get:
Which is identical to the form for dense angle encoding as seen in Eq. 3. When considering the multiqubit case, we must also incorporate the entanglement operator into our state. However, the dense angle encoding structure remains the same with changes in the phase value. A multiqubit formulation of QPMeL as a dense encoding would be:
| where, | |||
This is useful as it allows for QPMeL to leverage some of the theoretical properties of dense angle encoding. As shown by (LaRose and Coyle, 2020), dense angle encoding
The key contribution of QPMeL over standard angle encoding is coupling our classical encoder to work directly off this paradigm via the split angle embeddings. This creates an Intermediary Feature Space, which is curved rather than flat. This is enforced by the sigmoid function and the inherent normality constraints placed on and when they are encoded onto the Hilbert Space. This is shown in Fig. LABEL:fig:intermediate_space.
5.1.1. Bloch Sphere as a Spherical Embedding in QPMeL
The idea of spherical feature spaces in the context of metric learning has been explored by both (Wang et al., 2017) and (Zhang et al., 2020). However, while (Wang et al., 2017) had to implement a novel angular constraint for their triplet loss to create the angular constraints, the use of fidelity distance naturally enforces this constraint on the triplets.
A more interesting line of further study would be to look into applying the work done by (Zhang et al., 2020) to the QPMeL framework. Their team studied the effect of normalization on the gradients and found that proper normalization in the classical space can lead to better convergence. This is particularly interesting as the normalization of the qubit states is a natural part of the QPMeL framework. Their proposed modification to the triplet loss function normalized the embeddings before computing the distances:
The analog for QPMeL is:
In the QPMeL framework, our distance metric (fidelity) operates on quantum states which are normalized by nature. This would imply that it would inherit the desirable properties of the spherical embeddings. However, to prove that these properties hold for the gradients generated by parameterized quantum circuits would require further study and be beyond the scope of this work.
5.2. QPMeL as a Kernel Learner
As (Schuld, 2021) proved, the space of quantum models is mathematically identical to the reproducing kernel Hilbert space (RKHS) of Quantum Kernels. In this context, we can consider our classical head a kernel learner similar to the one proposed by (Le and Xie, 2018).
The proof and associated kernel structures can be seen below:
5.2.1. QPMeL Kernel Function
This is by design identical to the state fidelity due to the function of a kernel as a similarity metric. As ““ in the embedding from classical space, for QPMeL , we can define the embedding function as,
Therefore, the kernel function in the QPMeL framework is parameterized by defined as,
The QPMeL data encoding feature map was defined in Eqn. 2. This embedding function is significantly more complex than the embedding defined by a single rotational embedding considered by (Schuld, 2021). For the state of simplicity, let us only consider the rotational embedding components while ignoring the entanglement operator (as would be the case when ).
In order to better understand the kernel function, let us consider the case of a single qubit. As (Schuld, 2021) formulated the kernel function in terms of density matrices, I will be using the density matrice version of QPMeL.
The density matrix for our single qubit encoded state would be:
Therefore the inner product of two such states would yield our kernel function,
where,
This is quite similar to the kernel function of the linear model that (Schuld, 2021) uses. However, their paper used 1-D feature vectors alongside multiple parameters for the rotational gates forming the VQC. In contrast, QPMeL has 2-D feature vectors and only a single set of parameters for variational ZZ gates.
5.2.2. QPMeL as a Deep Embedding Kernel
As mentioned at the beginning of subsection. 5.2, we can consider the QPMeL as an extension of the Deep Embedding Kernel network introduced by (Le and Xie, 2018). This can be justified by looking at the structural similarities between the two frameworks as seen in Fig. LABEL:fig:QPMEL_DEK. We can further formalize the similarities between the two frameworks as follows:
The output of the Kernel network from (Le and Xie, 2018) is the probability that the inputs belong to the same class. In QPMeL the output of our training circuit is the fidelity between 2 encoded states which is also the probability that the inputs belong to the same class.
The embedding network in DEK(Le and Xie, 2018) creates optimal lower dimensional representations of the data based on the loss from the kernel network. In QPMeL , the classical head performs the same function.
Based on these parallels, we can define QPMeL as a Deep Embedding Quantum Kernel(DEQK) with the simplified kernel defined in Eq. 11.
@@ 6. Experimental Set-Up
6.1. MinMax Metric:
We propose the ”MinMax” metric to quantify our embedding performance. The metric is feature space agnostic as it directly operates on the pre-computed distances using the native space distance metric (MSE or fidelity).
However, as the metric trends are reversed for MSE and Fidelity, we present the results separately. Additionally, we use for Fidelity so that the sign implications of the metric are consistent. The metric computes the distance of the and , which corresponds to the worst case for both and represents it as a % . These are computed on the averages seen in the heatmap to avoid outliers.
AP
Worst case for Anchor-Positive distance.
The maximum distance between the same class.
The minimum distance between the same class.
AN
Worst case for Anchor-Negative distance.
The minimum distance between different classes.
The maximum distance between different classes.
@@ 6.1. MinMax Metric:
6.2. Classical Baselines:
6.3. Quantum Baselines:
6.4. Comparable Methods:
As mentioned in Section. 3, most of the comparable methods are limited to binary classification. Our proposed method
(Decision Boundry at 0.1)
(No Decision Boundry)
(No Decision Boundry)
(Decision Boundry at 1.5)
7. Results
(Perfect Separation)
(Smaller Perfect Separation)
(Multiple Outliers)
(4-6 pair is not separable)
Hilbert Compare.
7.1. Efficiency Analysis
When compared to the original QMeL Paper, we use 1/2 the number of gates and 1 layer circuit depth to achieve 3x better separation. The original QMeL framework (Lloyd et al., 2020) as explained in Section. 3 used an overlap loss that did not scale well to multi-class tasks. The modified version with our Fidelity Triplet Loss also failed to create a decision boundary as seen in Table.1 where it produces a negative difference. When we apply a more robust separation method for the pre-quantum step (QMeL+) the difference in Table.1 is positive implying that a decision boundary can be made but the magnitude is only 24% of implying that the clusters are close together.
In contrast, QPMeL produces a positive difference that is 91% of , showing 3x improvement over QMeL+. We see this trend reflected in Fig.6(a) and Fig.6(b) with the difference in magnitude between the diagonal and everywhere else. Additionally, looking at the parameters in Table 2 QPMeL utilizing 1/2 (9 vs 21) the number of gates, 1/2 (5 vs 11) the circuit depth and (11k vs 16k)20% fewer classical parameters compared to QMeL/QMeL+
@@ 7.2. Quantum loss functions may improve the learning of classical models
Look at Table 3, both sigmoid models produce negative differences which indicate the absence of a decision boundary. The QPMeL classical head has an identical structure to the Scaled Sigmoid model with only a difference in the loss function. However, it produces a positive difference of 88% of , implying that a strong decision boundary exists.
@@ 7.3. QPMeL strongly Couples Euclidean and Hilbert embedding seperation
Despite never being trained on any Euclidean loss function such as Mean Squared Error(MSE)-Triplet Loss, the QPMeL classical head produces a strong separation between classes as seen by the 88% of in Table 3. We can also see in Fig.8 that this separation is well distributed across all pairs.
@@ 7.4. QRC enables the classical head to learn decision boundaries in Hilbert Space
From the results in Table. 1, we can see that the No Residual model fails to learn a clear decision boundary. However, looking Fig.6(c), while the model fails to create a decision boundary, the issues are localized to specific pairs (ex. (8,2)), while all other regions remain well separated. We believe this is due to smaller gradients which require more training time and data to learn these corner cases. However, PQCs are known to have barren plateaus which can make training unstable.
In contrast, our model can learn a decision boundary for all classes. Fig.6(d), shows that even before applying our corrections the encoder struggles with a single pair. This implies that QRC is helping the classical model learn the corner cases faster. This is further supported by the fact that the QRC is only used during training and not during inference. This hints that the residual framework eases the task on the classical head allowing it to learn faster and more robustly.